# Estimation of AMOC transition probabilities using a machine learning based rare-event algorithm

### v1

@@ 4 Results: S-transitions
The S-transition case corresponds to a more complex problem for TAMS because the time scale of such a transition is much longer (about $1,000$ model years) and no score function can be derived from physical intuition. Castellana et al. (2019) also mentions this case but did not study it for a wide range of model parameters. For the S-transitions, we study $10$ values of $\overline{E_{a}}\in[0.09,0.345]$ and $f_{\sigma}\in[0.05,0.5]$, spanning the bistability regime. Here, as the equilibria are known, the prescribed score function (Baars et al., 2021) is:
@@
where |||||$\cdot$||| | ⋅ | | stands for the Euclidean 2-norm and $\eta=||\mathbf{x}^{U}-\mathbf{x}^{ON}||/||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||$ corresponds to the distance between the on-state $\mathbf{x}^{ON}$ and the unstable steady state $\mathbf{x}^{U}$, normalised by the distance between $\mathbf{x}^{ON}$ and the steady off state $\mathbf{x}^{OFF}$. This score function makes sense from a mathematical point of view (the transition between two steady states should go through a saddle-node) but does not allow any physical interpretation. Moreover, it may be strongly biased when $f_{\sigma}$ is large or when the system has multiple unstable steady states. In this case, it is particularly interesting to estimate the committor function via the RC method. We keep the same parameters for TAMS, except for $T_{\mathrm{max}}=1,000$ years, and still call TAMS-$S$ the combination of TAMS with $\Phi_{S}^{S}$ and TAMS-$R$ the combination of TAMS with $\Phi_{R}^{S}$.
TAMS-$R$ is still initialized using an ensemble of $1,000$ pre-generated trajectories with $(\overline{E_{a}}=0.345\mathrm{Sv},f_{\sigma}=0.5)$. But in this case, $\alpha_{S}^{S}=0.078$ so to prevent any bias, we use $500$ trajectories that don’t contain a transition (obtained by direct numerical simulation) $500$ trajectories containing a transition (obtained with TAMS-$S$).
Figure 6: a) Mean probability to transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$ within $1000$ years, over the whole range of values of $(\overline{E_{a}},f_{\sigma})$. This probability was averaged over $30$ runs of TAMS-$S$.
@@
Figure 6a shows the diagram of the mean transition probabilities from the AMOC on state to the AMOC off state, computed with $\Phi_{S}^{S}$ and used as reference. Fig. 6b presents the consistency score $\mathcal{C}$ between TAMS-$S$ and TAMS-$R$. Figure 6a shows that, as expected, the smaller the freshwater forcing $\overline{E_{a}}$, the smaller the transition probabilities, as for the F-transitions. However, now a large noise amplitude $f_{\sigma}$ tends to prevent the transition to $\mathbf{x}^{OFF}$. This can be understood by considering the basin of attraction of $\mathbf{x}^{ON}$. Since this basin is very narrow, it only requires a small noise to exit it and be attracted to $\mathbf{x}^{OFF}$. In this setting, a small noise already allows the system to switch basins of attraction and does not disturb it from its natural deterministic drift toward $\mathbf{x}^{OFF}$. On the other hand, a large noise strongly disturbs the system and may force it in directions that do not favor a transition to $\mathbf{x}^{OFF}$, thus lowering the probability of reaching it within $1,000$ years.
@@
The first element to notice in Fig. 6b is that it contains fewer colored grid cells than in 6a. This is due to the fact that the RC was unable to estimate a nonzero probability of collapse when $\alpha_{S}^{S}<10^{-6}$. This failure of the RC for very low transition probabilities may be explained by a less efficient transfer learning process than in the case of the transient collapse. Indeed, the gap between successive sampled parameters $(\overline{E_{a}},f_{\sigma})$ is $10$ times larger than in the previous case, making it more difficult for the RC to adapt to new dynamics. However, this also means that even under these poor learning conditions, the RC can still estimate transition probabilities down to $10^{-6}$ for a wide range of parameters $(\overline{E_{a}},f_{\sigma})$. Namely, the predicted probability $\alpha_{R}^{S}$ is consistent with $\alpha_{S}^{S}$ (so $\mathcal{C}<1$, shown in blue) for $83\%$ of the parameter couples. In all cases where this is not verified (except for one), the inconsistently estimated transition probabilities are the lowest for a given value of $f_{\sigma}$.
@@
Figure 7: (a) Ratio of the number of timesteps computed by TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ require fewer timesteps, blue values mean the opposite.
@@
(b) Ratio of the number of iterations of TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ requires fewer iterations, blue values mean the opposite.
@@
(c) Ratio of the variance in probability estimates obtained with TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ produces a larger variance, blue values mean the opposite. In all panels, the gray color indicates that $\alpha_{S}^{S}$ or $\alpha_{R}^{S}$ is below the $10^{-9}$ cut-off threshold.
@@
Figure 7a compares how many time steps are computed by both methods. Once again, for high probabilities, both methods are, as expected, equivalent because TAMS converges very fast. When they are not equivalent, TAMS-$R$ almost always leads to the computation of more time steps, which is shown in blue in Fig. 7a. The difference is even quite large: in $50\%$ of the cases, TAMS-$R$ leads to the computation of more than $40\%$ more time steps compared to TAMS-$S$. However, if we only consider low probabilities $\alpha_{S}^{S}<10^{-3}$, TAMS-$R$ computes more than $40\%$ more time steps compared to TAMS-$S$ for $40\%$ of the parameter values, and this difference in the calculated time steps is less than $30\%$ of the time steps of TAMS-$S$ for $30\%$ of the parameter values. Therefore, the number of additional time steps computed by TAMS-$R$ compared to TAMS-$S$ stabilizes for lower probabilities.
@@
However, this difference in the number of calculated time steps hides the fact that TAMS-$R$ converges faster than TAMS-$S$. The number of TAMS iterations for both methods is presented in Fig. 7b, and for $90\%$ of the parameters, the number of iterations of TAMS-$R$ is smaller than that of TAMS-$S$. Even when it is larger, the difference is always less than $3\%$ of the number of iterations of TAMS-$R$ (except in two cases where this difference reaches up to $10\%$ of the number of iterations of TAMS-$R$, for probabilities lower than $10^{-4}$). For half of the parameter values, this difference corresponds to more than $15$ more iterations performed by TAMS-$S$ but can reach more than $100$ more iterations in $20\%$ of the parameter values. That TAMS-$R$ performs fewer iterations to converge is consistent with the fact that it computes more time steps: if TAMS converges earlier, it means that it was easier for the trajectories to reach a transition, which implies that they must have been systematically branched at an earlier time. More importantly, a smaller number of iterations also suggests a better score function, since detection of promising and wrong trajectories occurs earlier within the fixed time interval $[0,T_{\mathrm{max}}]$.
@@
Figure 7c represents the variance around the probability estimate for both methods. As is already the case with the other panels, both methods are similar for large transition probabilities: the difference in the values of $(\sigma^{2})^{S}$ is less than $5\%$ of $(\sigma^{2})_{S}^{S}$ only for transition probabilities larger than $0.5$. $(\sigma^{2})_{R}^{S}$ is only smaller than $(\sigma^{2})_{S}^{S}$ for transition probabilities $p\geq 0.08$. Otherwise, TAMS-$S$ systematically yields a smaller variance, with a difference that can be quite large: $\alpha_{R}^{S}$ is more than $50\%$ larger than $\alpha_{S}^{S}$ for half of the parameter values. As is also the case in the other panels of Fig. 7, the smaller the transition probability, the larger this discrepancy and, here, the larger $(\sigma^{2})_{R}^{S}$ compared to $(\sigma^{2})_{S}^{S}$. As we move towards the left of the diagram, the difficulty of the transfer learning does not allow the RC to adapt to new dynamics, and its estimates are increasingly scattered. For even lower probabilities, TAMS-$R$ cannot even converge in many of the $30$ runs but gives a correct probability estimate in others, leading to a very large variance. This also explains why, to the very left of the diagram, there are some cases where $(\sigma^{2})_{R}^{S}<(\sigma^{2})_{S}^{S}$: in this case, there are so many TAMS runs that did not converge and returned a probability of 00 that the prediction of TAMS-$R$ becomes quite consistent. Unlike the variances presented in Fig. 3, the variances shown in Fig. 7c are computed over $30$ runs of TAMS-$R$, each using an independently trained RC.
@@ 5 Interpretability of the RC
An important advantage of using RCs is that they can be interpreted. Each column of $\mathbf{W}_{\mathrm{out}}$ (see Appendix C) corresponds to a combination of powers of $S_{n},S_{ts},D$ and $t$. We can thus read in this matrix the terms that contribute the most to the estimated committor.
@@
Figure 8: Each panel shows, for a fixed $f_{\sigma}$, the terms with largest magnitude in the matrix $W_{\mathrm{out}}$ of the RCs. From the $30$ RCs trained in section 3.2, we extract the $10$ terms of $W_{\mathrm{out}}$ with largest magnitude. Their magnitude is then normalized to that of the largest term. We then only keep the terms that are most significant for more than $20$ out of $30$ RCs. Each curve correspond to a value of the parameter $\overline{E_{a}}$ and they show the average magnitude of each term across the $30$ RCs and the errorbars show the corresponding standard deviations.
First, we focus on $\Phi_{R}^{F}$. In the case of F-transitions, we have run TAMS $30$ times for $30$ different RCs at $3$ values of $f_{\sigma}$ (see Sect. 3.2). For each of the RCs, we select the $10$ values of $W_{\mathrm{out}}$ that have the largest amplitude (in absolute value), and we only keep the terms that are common in more than $20$ RCs. Their average magnitude (across the $30$ RCs) is normalized so that the largest is equal to $1$ in absolute value. Figure 8 presents these magnitudes with their standard deviation across the $30$ RCs. For each value of $f_{\sigma}$, if the most significant terms are similar for several values of $\overline{E_{a}}$, they are plotted together on the corresponding panel.
@@
From Fig. 8, we see that for all values of $f_{\sigma}$, the most important terms of all RCs are very consistent: few error bars on the magnitude of every term are visible. This means that all RCs can extract a meaningful structure from the data. Furthermore, for each $f_{\sigma}$, the structure of the matrix $W_{\mathrm{out}}$ also converges as the transition probabilities become very small, as shown by the overlap of the curves obtained for several values of $\overline{E_{a}}$.
@@
These curves also allow for an analytical reconstruction of a suitable score function. For $f_{\sigma}=0.25$ and $f_{\sigma}=0.5$, $3$ interesting terms are highlighted: $S_{n}S_{ts}D^{2}$ (with a magnitude of $-1$) and $S_{n}^{2}D^{2}$ and $S_{ts}^{2}D^{2}$ (with a magnitude of about $0.5$). This naturally leads to a score function $\Phi_{R}^{F}\propto(S_{n}-S_{ts})^{2}D^{2}/2$. In both cases, we can also add the terms $(S_{ts}-S_{n})S_{ts}t^{2}/2$ and $(S_{n}^{2}+S_{ts}^{2}/2-S_{n}S_{ts})Dt/2$. These terms are not exactly contained in the coefficient of $W_{\mathrm{out}}$ because the coefficients of this matrix are not exactly $1$, $0.5$, or $0.25$. These analytical expressions give, nonetheless, a first estimate of an analytical formula for $\Phi_{R}^{F}$, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$:
@@
for certain ${\cal O}(1)$ coefficients $a_{i},\ i=1,2,3$, which have appropriate dimensions to make the expression consistent. This expression does not have an ”equal” sign because it is rather a proportionality relation, due to the normalization of $W_{\mathrm{out}}$. $\Phi_{R}^{F}$ can be compared to the structure of $\Phi_{S}^{F}$, driven by the term $(S_{n}-S_{ts})D^{2}$. Here, this precise term cannot be retrieved because the nonlinear part of the feature vector contains only monomials of degree $4$ (see Appendix C). The original implementation of the RC (Gauthier et al., 2021), which we followed, imposes a fixed choice of degree and we found $4$ to work best (Jacques-Dumas et al., 2023). Although $\Phi_{S}^{F}$ cannot be exactly reproduced, the RCs retrieve a similar dependence on $D^{2}$ and $(S_{n}-S_{ts})$. Most importantly, $\Phi_{R}^{F}$ involves time, which is what we should expect from a score function, since the transition must occur before $T_{\mathrm{max}}$. In particular, the third term of Eq. 6 is very interesting, as it combines $D$, $t$, and a term that looks like $(S_{n}-S_{ts})^{2}$, which would not have been possible if the degree of the monomials in the feature vector had been smaller.
Finally, it is striking to notice that the expressions for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ are very close. This is due in part to transfer learning, but, considering that the RCs are constantly independently retrained, the conservation of this structure across different values of $\overline{E_{a}}$ and $f_{\sigma}$ shows the robustness of this method. However, the RC is still able to adapt, as shown in the third panel of Fig. 8 by the very different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$. In this case, two monomials stand out: $S_{n}^{2}D^{2}$ and $-S_{ts}^{2}D^{2}$. This suggests a score function $\Phi_{R}^{F}(S_{n},S_{ts},D,t)\propto(S_{n}-S_{ts})(S_{n}+S_{ts})D^{2}$. Other significant terms in these curves are $-(S_{n}^{2}-S_{ts}^{2})Dt/4$ and $(S_{n}^{2}-S_{ts}^{2}/2)t^{2}/2$, leading to a new analytical expression of the score function obtained by the RCs:
@@
for certain ${\cal O}(1)$ coefficients $b_{i},\ i=1,2,3$. The estimated score function for $f_{\sigma}=0.05$ is different from the score obtained for much larger values of the noise amplitude, since a different noise induces different dynamics in the model. Thus, a data-driven score function is expected to depend on the model parameters. If we were able to solve the backward Kolmogorov equation (see Appendix B), the exact committor would depend on $f_{\sigma}$, which would make the structure of the score function evolve with $f_{\sigma}$. However, here such an analytical dependence on $f_{\sigma}$ cannot be retrieved and only its effect can be observed. Finally, despite the different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$, its dependence on $(S_{n}-S_{ts})$ is conserved, although always multiplied by $(S_{n}+S_{ts})$, stressing more the importance of the salt-advection feedback. The dependence on three terms driven by $D^{2}$, $Dt$ and $t^{2}$ is also conserved.
Such an analysis may be very useful in more complex cases and larger-dimensional models to understand the possible structure of a good score function. We test this point by interpreting the output of $W_{\mathrm{out}}$ when trained to study the transition from the AMOC steady on state to the AMOC steady off-state (studied in Sect. 4).
@@
Fig. 9 shows the same type of curves as Fig. 8, obtained from the matrix $W_{\mathrm{out}}$ of $30$ RCs trained on an S-transition of the AMOC, for two values of $f_{\sigma}$: $0.5$ and $0.1$. For each value of $f_{\sigma}$, several curves are presented, corresponding to probability estimates $\alpha_{R}^{S}$ between $0.002$ and $0.1$. For both values of $f_{\sigma}$, all curves are very consistent, showing the ability of the RCs to capture some structure of the phase space and to converge, although they are constantly independently retrained. However, in both panels, the curves corresponding to the lowest transition probabilities have large error bars, showing the limits of transfer learning with a large difference between consecutive values of $\overline{E_{a}}$. In this case, the analytical formulas for the committor that can be extracted from these curves are:
@@
for again specific coefficients $c_{i}$ and $d_{i}$. These expressions are difficult to relate to $\Phi_{S}^{S}$ because the latter relied purely on Euclidean distances in phase space rather than on some specific physical variables.

### v3

@@ 4 Results: S-transitions
The S-transition case corresponds to a more complex problem for TAMS because the time scale of such a transition is much longer (about $1,000$ model years) and no score function can be derived from physical intuition. Castellana et al. (2019) also mentions this case but did not study it for a wide range of model parameters. For the S-transitions, we consider $10$ equally spaced values of $\overline{E_{a}}\in[0.09,0.345]\mathrm{\ Sv}$ (spanning the bistability regime) and $f_{\sigma}\in[0.05,0.5]$. Here, as the equilibria are known, the prescribed score function (Baars et al., 2021; Castellana et al., 2019) is:
@@
where |||||$\cdot$||| | ⋅ | | stands for the Euclidean 2-norm and $\eta=||\mathbf{x}^{U}-\mathbf{x}^{ON}||/||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||$ corresponds to the distance between the on-state $\mathbf{x}^{ON}$ and the unstable steady state (saddle-node) $\mathbf{x}^{U}$, normalized by the distance between $\mathbf{x}^{ON}$ and the steady off state $\mathbf{x}^{OFF}$. This score function makes sense from a mathematical point of view (the transition between two steady states should go through a saddle-node) but does not allow any physical interpretation. Moreover, it may be strongly biased when $f_{\sigma}$ is large or when the system has multiple unstable steady states. In this case, it is particularly interesting to estimate the committor function with the RC. We keep the same parameters for TAMS, except for $T_{\mathrm{max}}=1,000$ years, and still call TAMS-$S$ the combination of TAMS with $\Phi_{S}^{S}$ and TAMS-$R$ the combination of TAMS with $\Phi_{R}^{S}$.
TAMS-$R$ is initialized using an ensemble of $1,000$ pre-generated trajectories with $(\overline{E_{a}}=0.345\mathrm{\ Sv},f_{\sigma}=0.5)$. But in this case, $\alpha_{S}^{S}=0.078$ so to prevent any bias, $500$ trajectories don’t contain a transition (obtained by direct numerical simulation) and $500$ trajectories contain one (obtained with TAMS-$S$). Otherwise, due to the rarity of transitions, $1000$ trajectories would contain too few of them and such unbalanced training dataset would hinder performance.
Figure 6: a) Mean estimated probability to transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$ within $1000$ years, over the whole range of values of $(\overline{E_{a}},f_{\sigma})$. This probability was averaged over $30$ runs of TAMS-$R$.
@@
Figure 6a shows the diagram of the mean transition probabilities from the AMOC on state to the AMOC off state, computed with $\Phi_{R}^{S}$. Fig. 6b presents the consistency score $\mathcal{C}$ between TAMS-$S$ and TAMS-$R$. Figure 6a shows that, as expected, the smaller the freshwater forcing $\overline{E_{a}}$, the smaller the transition probabilities, as for the F-transitions. However, now, a large noise amplitude $f_{\sigma}$ tends to prevent the transition to $\mathbf{x}^{OFF}$. This can be understood by considering the basin of attraction of $\mathbf{x}^{ON}$. Since this basin is very narrow, it only requires a small noise to exit it and be attracted to $\mathbf{x}^{OFF}$. In this setting, a small noise already allows the system to switch basins of attraction and does not disturb it from its natural deterministic drift toward $\mathbf{x}^{OFF}$. On the other hand, a large noise strongly disturbs the system and may force it in directions that do not favor a transition to $\mathbf{x}^{OFF}$, thus lowering the probability of reaching it within $1,000$ years. The cut-off to compute transition probabilities with TAMS was here as well set to $10^{-9}$ but TAMS-$R$ reaches such a low probability only for $(\overline{E_{a}}=0.15\ \mathrm{Sv},f_{\sigma}=0.25)$. In all other cases, it fails to compute transition probabilities lower than $10^{-6}$. This failure of the RC for very low transition probabilities may be explained by a less efficient transfer learning process than in the case of the transient collapse. Indeed, the gap between successive sampled parameters $(\overline{E_{a}},f_{\sigma})$ is $10$ times larger than in the previous case, making it more difficult for the RC to adapt to new dynamics.
@@
However, this also means that even under these poor learning conditions, the RC can still estimate transition probabilities down to $10^{-6}$ for a wide range of parameters $(\overline{E_{a}},f_{\sigma})$. Fig. 6b demonstrates that the predicted probability $\alpha_{R}^{S}$ is consistent with $\alpha_{S}^{S}$ (so $\mathcal{C}<1$, shown in blue) for $83\%$ of the parameter couples. In all cases where this is not verified (except for one), the inconsistently estimated transition probabilities are the lowest for a given value of $f_{\sigma}$.
@@
Figure 7: (a) Ratio of the number of timesteps computed by TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$R$ require fewer timesteps, blue values mean the opposite. In all panels, the gray color indicates that $\alpha_{S}^{S}$ or $\alpha_{R}^{S}$ is below the $10^{-9}$ cut-off threshold.
@@
(b) Ratio of the number of iterations of TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$R$ requires fewer iterations, blue values mean the opposite.
@@
(c) Ratio of the variance in probability estimates obtained with TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ produces a larger variance, blue values mean the opposite.
@@
Figure 7a compares how many time steps are computed by both methods. Since each iteration of TAMS consists in integrating trajectories, i.e. simulating many time steps, we summed them up during the whole TAMS process to obtain an indicator of its computational cost. For high probabilities, both methods are, as expected, equivalent because TAMS converges very fast. When they are not equivalent, TAMS-$R$ almost always leads to the computation of more time steps, which is shown in blue in Fig. 7a. The difference is even quite large: in $50\%$ of the cases, TAMS-$R$ leads to the computation of more than $40\%$ more time steps compared to TAMS-$S$.
@@
However, the smaller number of time steps computed by TAMS-$S$ hides the fact that the latter takes more iterations to converge than TAMS-$R$. The number of TAMS iterations for both methods is presented in Fig. 7b, and for $90\%$ of the parameters, the number of iterations of TAMS-$R$ is smaller than that of TAMS-$S$. Even when it is larger, the difference is always less than $3\%$ of the number of iterations of TAMS-$R$ (except in two cases where this difference reaches up to $10\%$ of the number of iterations of TAMS-$R$, for probabilities lower than $10^{-4}$). For half of the parameter values, this difference corresponds to more than $15$ more iterations performed by TAMS-$S$ but can reach more than $100$ more iterations in $20\%$ of the parameter values. The fewer iterations of TAMS-$R$ are consistent with its larger number of computed time steps: to converge in fewer iterations, trajectories must be branched earlier (within $[0,T_{\mathrm{max}}]$) so their resimulation involves the computation of more time steps. More importantly, this also suggests a better score function, since detection of promising and wrong trajectories occurs earlier within the time interval $[0,T_{\mathrm{max}}]$.
@@
Figure 7c represents the variance around the probability estimates for both methods. As in the other panels, both methods are similar for large transition probabilities: the difference in the values of $(\sigma^{2})_{R}^{S}$ and $(\sigma^{2})_{S}^{S}$ is less than $5\%$ of $(\sigma^{2})_{S}^{S}$ only for transition probabilities larger than $0.5$. $(\sigma^{2})_{R}^{S}$ is only smaller than $(\sigma^{2})_{S}^{S}$ for transition probabilities $\alpha_{R}^{S}\geq 0.08$. Otherwise, TAMS-$S$ systematically yields a smaller variance, with a difference that can be quite large: $\alpha_{R}^{S}$ is more than $50\%$ larger than $\alpha_{S}^{S}$ for half of the parameter values. As is also the case in the other panels of Fig. 7, the smaller the transition probability, the larger this discrepancy and, here, the larger $(\sigma^{2})_{R}^{S}$ compared to $(\sigma^{2})_{S}^{S}$. As we move towards the left of the diagram, the difficulty of the transfer learning does not allow the RC to adapt to new dynamics, and its estimates are increasingly scattered. For even lower probabilities, TAMS-$R$ cannot even converge in many of the $30$ runs but gives a correct probability estimate in others, yielding a very large variance. This also explains why, to the very left of the diagram, there are some cases where $(\sigma^{2})_{R}^{S}<(\sigma^{2})_{S}^{S}$: in this case, there are so many TAMS runs that did not converge and returned a probability of 00 that the prediction of TAMS-$R$ becomes quite consistent. Note that the variances presented in Fig. 3 were computed by running TAMS $30$ times for each of the $30$ independently trained RCs whereas the variance plotted in Fig. 7c is directly computed from $30$ runs of TAMS, without freezing the RCs.
@@ 5 Interpretability of the RC
Figure 8: Each panel shows, for a fixed $f_{\sigma}$, the terms with largest magnitude in the matrix $W_{\mathrm{out}}$ of the RCs. From the $30$ RCs trained in Sect.3b, we extract the $10$ terms of $W_{\mathrm{out}}$ with largest magnitude. Their magnitude is then normalized to that of the largest term and we only keep the terms that are most significant for more than $20$ out of $30$ RCs. Each curve corresponds to a value of the parameter $\overline{E_{a}}$ (the legend gives the corresponding transition probability computed with TAMS-$R$) and they show the average magnitude of each term across the $30$ RCs and the errorbars are the corresponding standard deviations.
An important advantage of RCs is that we can read which term of their output matrix $W_{\mathrm{out}}$ contributed most to the committor estimation. In the case of F-transitions, we consider the $30$ runs of TAMS-$R$ performed at $f_{\sigma}=0.5,0.25$ and $0.05$. For each of the $30$ RCs, we select the $10$ largest terms of $W_{\mathrm{out}}$ in absolute value and normalize them, so that the largest coefficient in each RC has a magnitude of $1$. Then, we only keep the terms that are common in more than $20$ RCs (the results are robust when varying this number between $15$ and $25$), to compare how the most significant terms are ranked by each RC. Figure 8 presents these relative magnitudes with their standard deviation across the $30$ RCs. Each curve corresponds to a value of $\overline{E_{a}}$ (or equivalently to a grid cell in Fig. 2a). For a given value of $f_{\sigma}$, they were sampled in the following way: starting from the smallest value of $\overline{E_{a}}$ such that $\alpha_{R}^{F}\geq 10^{-9}$, we retained one value of $\overline{E_{a}}$ out of $10$ such that $\alpha_{R}^{F}\leq 0.5$. By imposing a certain interval on the transition probabilities, we make sure that all the presented curves roughly correspond to similar dynamics of the model despite the different values of $(\overline{E_{a}},f_{\sigma})$.
@@
Figure 8 shows that, for all values of $f_{\sigma}$, the most important terms of all RCs are consistent: few error bars on the magnitude of every term are visible. All RCs thus extract a meaningful structure from the data and they all agree on the ranking of the most significant terms. Furthermore, for each $f_{\sigma}$, the structure of the matrix $W_{\mathrm{out}}$ converges as the transition probabilities become very small, as shown by the overlap of the curves obtained for several values of $\overline{E_{a}}$.
@@
These curves also provide an analytical expression of the learned score function. For $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, $3$ interesting terms are highlighted: $S_{n}S_{ts}D^{2}$ (with a magnitude of $-1$) and $S_{n}^{2}D^{2}$ and $S_{ts}^{2}D^{2}$ (with a magnitude of about $0.5$). This naturally leads to a score function $\Phi_{R}^{F}\propto(S_{n}-S_{ts})^{2}D^{2}/2$. In both cases, we can also add the terms $(S_{ts}-S_{n})S_{ts}t^{2}/2$ and $(S_{n}^{2}+S_{ts}^{2}/2-S_{n}S_{ts})Dt/2$ although these terms are not exactly contained in $W_{\mathrm{out}}$ because the coefficients of this matrix are not exactly $1$, $0.5$, or $0.25$. We still obtain a first estimate of an analytical result for $\Phi_{R}^{F}$, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$:
@@
for certain ${\cal O}(1)$ coefficients $a_{i},\ i=1,2,3$ with appropriate dimensions to make the expression consistent.
Due to the normalization of $W_{\mathrm{out}}$, this expression is only a proportionality relation. Moreover, the used implementation of the RC by Gauthier et al. (2021) imposes a fixed choice of degree of the terms in the nonlinear part of the feature vector and we found $p=4$ to work best (Jacques-Dumas et al., 2023). Although $\Phi_{S}^{F}\propto q_{N}\propto(S_{n}-S_{ts})D^{2}$ cannot be reproduced (because $q_{N}$ is of degree $3$), we can compare this expression to that of $\Phi_{R}^{F}$. The RCs retrieve a similar dependence on $D^{2}$ and $(S_{n}-S_{ts})$. This salinity difference drives the density gradient between the boxes n and ts, which, in turn, controls in this model the AMOC strength $q_{N}$ (Cimatoribus et al., 2014), implying that the salt advection drives the AMOC strength and that the salt advection feedback is key in the AMOC collapse. The presence of $(S_{n}-S_{ts})$ in all terms of $\Phi_{R}^{F}$ shows that the RC is able to capture the importance of this mechanism. $\Phi_{R}^{F}$ also involves time, which was to be expected from the score function. Indeed, TAMS returns a transition probability conditioned on a time horizon $T_{\mathrm{max}}$ so the score function is ideally time-dependent. In particular, the third term of Eq. 5 combines $D$, $t$, and a term looking like $(S_{n}-S_{ts})^{2}$, which would not be possible if the degree of the monomials was $p<4$.
Finally, the expressions for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ are strikingly close. This is partly due to transfer learning, but, considering that the RCs are constantly independently updated, the conservation of this structure across different values of $\overline{E_{a}}$ and $f_{\sigma}$ shows the robustness of this method. However, the different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$ (third panel of Fig. 8) shows that the RCs are still able to adapt. In this case, two monomials stand out: $S_{n}^{2}D^{2}$ and $-S_{ts}^{2}D^{2}$. This suggests a score function $\Phi_{R}^{F}(S_{n},S_{ts},D,t)\propto(S_{n}-S_{ts})(S_{n}+S_{ts})D^{2}$. Other significant terms in these curves are $-(S_{n}^{2}-S_{ts}^{2})Dt/4$ and $(S_{n}^{2}-S_{ts}^{2}/2)t^{2}/2$, leading to a new analytical expression of the score function obtained by the RCs:
@@
for certain ${\cal O}(1)$ coefficients $b_{i},\ i=1,2,3$. The function $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$ is different from the score function obtained for larger values of $f_{\sigma}$, which is not surprising as a data-driven score function is expected to depend on the model parameters. Moreover, $f_{\sigma}$ already appears in the equations of the committor (Lucente et al., 2022), so this function is expected to depend on it. Finally, despite the different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$, its dependence on $(S_{n}-S_{ts})$ is conserved, although always multiplied by $(S_{n}+S_{ts})$, stressing more the importance of the salt-advection feedback. The dependence on three terms driven by $D^{2}$, $Dt$ and $t^{2}$ is also conserved.
Such an analysis may be very useful in more complex cases and larger-dimensional models to understand the possible structure of a good score function. We test this point by interpreting the output of $W_{\mathrm{out}}$ in the case of the AMOC S-transitions (see Sect. 4).
@@
Fig. 9 shows the same type of curves as Fig. 8, obtained from the matrix $W_{\mathrm{out}}$ of $30$ RCs trained on an S-transition of the AMOC, for two values of $f_{\sigma}$: $0.5$ and $0.1$. Each curve corresponds to a grid cell in Fig. 6. For each value of $f_{\sigma}$, we display all curves for which the probability estimates $\alpha_{R}^{S}$ lies between $0.002$ and $0.1$. This corresponds to only two curves in panel b), because the jumps in transition probabilities between consecutive samples of $\overline{E_{a}}$ is larger for $f_{\sigma}=0.1$ than for $f_{\sigma}=0.5$ (as can be observed in Fig. 6a). For both values of $f_{\sigma}$, all curves are very consistent, showing the ability of the RCs to capture some structure of the phase space and to converge, although they are constantly independently retrained. For $f_{\sigma}=0.5$, the curve corresponding to the largest $\alpha_{R}^{S}$ deviates from the others (on the right of panel a)), although it has large error bars. In both panels, the curves corresponding to the lowest transition probabilities also have large error bars, showing the limits of transfer learning with a large difference between consecutive samples of $\overline{E_{a}}$. In this case, the analytical results for the committor that can be extracted from these curves are:
@@
for again specific ${\cal O}(1)$ coefficients $c_{i}$ and $d_{i}$. These expressions are difficult to relate to $\Phi_{S}^{S}$ because the latter only relies on Euclidean distances in phase space rather than specific physical variables.